OpenStudy (anonymous):

Solve: (x-2)^.66 = 49

OpenStudy (anonymous):

is that raised to 0.66?

OpenStudy (anonymous):

(x-2)^2/3 = 49

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cuberoot((x-2)^2) = 49

OpenStudy (anonymous):

can you answer that now?

OpenStudy (anonymous):

I'm not that good at math so I'm going to say no. haha

OpenStudy (anonymous):

\[\sqrt[3]{(x-2)^2} = 49\]

OpenStudy (anonymous):

to answr this, you must try to look on how you will be able to make the left hand side look similar to the right hand side

OpenStudy (anonymous):

clue: \[\sqrt[3]{(x-2)^2} \] is also equal to

OpenStudy (anonymous):

\[(\sqrt[3]{x-2})^2\]

OpenStudy (anonymous):

see something to work with?

OpenStudy (anonymous):

dude, lol I am just not even understanding this right now. My math class starts in like 15 minutes. Is there an easier way to explain?

OpenStudy (anonymous):

\[(\sqrt[3]{x-2})^2 = 49\]

OpenStudy (anonymous):

49 is also equal to???

OpenStudy (anonymous):

sorry, i can't say the answer directly. rules. hehhee

OpenStudy (anonymous):

try manipulating 49

OpenStudy (anonymous):

to MAKE IT LOOK SIMILAR TO THE LEFT HAND SIDE

OpenStudy (anonymous):

Hey, don't yell. Haha I understand tho. I really appreciate your help! Let's be friend ;)

OpenStudy (anonymous):

I'm not yelling. I just wanted to emphasis that statement

OpenStudy (anonymous):

in what ways can you make 49 look somewhat simalar to the left hand side?

OpenStudy (anonymous):

I know, I'm just trolling you lol

OpenStudy (anonymous):

But I really do need the help so yeah.

OpenStudy (anonymous):

we all know that 49 is a perfect square, right?

OpenStudy (anonymous):

yeah... 7

OpenStudy (anonymous):

so??? since there's \[(\sqrt[3]{x-2})2\]

OpenStudy (anonymous):

and there's 7^2

OpenStudy (anonymous):

what will you do now?

OpenStudy (anonymous):

if the exponents of 2 terms in an equality are the same, then they are equal to each other. You can just remove the ^2

OpenStudy (anonymous):

that leaves you where?

OpenStudy (anonymous):

so just 7? Man, this would be easier if I was talking to you face to face haha

OpenStudy (anonymous):

yes and how bout the left side?

OpenStudy (anonymous):

you have just removed the right side's exponent only

OpenStudy (anonymous):

you have just removed the right side's exponent only \[\sqrt[3]{x-2} = 7\]

OpenStudy (anonymous):

Are you getting this?

OpenStudy (anonymous):

Kind of... I'm following you a bit. Welp, math class is about to start so, lol I guess I'll keep in touch?

OpenStudy (anonymous):

just make sure to isolate x from all the constants